OpenStudy (anonymous):

write a polynomial function of least degree that has real coefficents, the given zeros, and a leading coefficent of 1 2,-2,i,-i

OpenStudy (anonymous):

see if x=2 be a root then x-2 is one of the factor

OpenStudy (anonymous):

To wirte a polynomial with give zeros you do (x-a)(x-b)(x-c)(x-d) where a b c and d are the zeros by the defiition of the linear factorizartion theorme

OpenStudy (anonymous):

Its already stated that they are zeros @matricked

OpenStudy (anonymous):

(x+2)(x-2)(x+i)(x-i)=(x^2-4)(x^2+1)= x^4-3x^2 -4